An airplane travels 400 miles against the wind in 5 hours, and makes the return trip with the same wind in 2hours. Find the speed of the wind.
Answer by stanbon(75887)

An airplane travels 400 miles against the wind in 5 hours, and makes the return trip with the same wind in 2hours. Find the speed of the wind.
-------------------------
Against the wind DATA:
distance = 400 miles ; time = 5 hrs ; rate = 400/5 = 80 mph
--------------------------------------------------------------
With the wind DATA:
distance = 400 miles ; time = 2 hrs ; rate = 400/2 = 200 mph
----------------------------------------------------------------
Equations:
Let "p" be speed of the plane in still air
Let "w" be speed of the wind
---------------------------------
p + w = 200
P - w = 80
---------------------------------
Add the two equations to solve for "p":
2p = 280
p = 140 mph (speed of the plane in still air)
---Substitute into P+w = 200 to solve for "w"
140 + w = 200
w = 60 mph (speed of the wind)
======================================
Cheers,
Stan H.